Reports 1-1 of 1 Clear search Modify search
Injection system (Laser power stabilization)
Bersanetti, Chaibi, chiummo, Daumas, Magazzu, Troiano - 22:57 Sunday 19 March 2023 (59325) Print this report
PSL power rebalance

This afternoon the rebalance of PSL power discarding has been performed by lowering the power sent to the PMC and increasing the transmission of IPC1 in order to keep the same power transmitted by IMC.

The power transmitted by the PMC has been reduced by roughly 18% by acting on the HWP/polarizers just upstream of the PSTAB AOM. We measured the beam profile coming out of the PMC before and after the power variation (results in a separate entry), and realigned the PMC by suing the steering mirrors after the AOM and before the PMC. This had the effect of improving the mismatch to the IMC cavity while delivering the same power from IMC.

Also, the reduced power sent into the AOM seemed to be beneficial to the PSTAB loop. We maanged to engage the PSTAB loop without seeing the glitches which wwere there during the past days - no gain tuning was deemed to be necessary at this moment.

Unfortunately, the drawback was that the IMC linear alignment works for the moment in drift control, a check of the working point and possibly the alignment matrix will be done as soon as possible. For the time being the engagement of full bandwidth linear alignment has been left commented in metatron.

The interferometer was locked up to carm_null_1f after the work, and left to the TCS measurements.

 

 

Images attached to this report
Comments to this report:
chiummo - 10:10 Monday 20 March 2023 (59331) Print this report

PSTAb behavior during last night confirmed that the situation has been improved after the reduction of input power in the AOM before the PMC. To be futher investigated.

Images attached to this comment
chiummo - 11:49 Monday 20 March 2023 (59333) Print this report

Comparing the mismatch towards the IMC in the two different configurations: purple is with max PMC transmission, blue with reduced power sent to PMC (current), comparable IMC_TRA power.

The IMC_REFL_DC channel is reported after offset substraction (V0 = 0.34V), divided by the EIB_Pout power and normalized to 1 when IMC is unlocked. IMC mismatch was around 16%, is now around 11%.

Images attached to this comment
Search Help
×

Warning

×